# Improving Detection in Aerial Images by Capturing Inter-Object Relationships

### v1

@@ 4.3 Comparison with Baselines
Transformer Relations ($P$) Ada. Weight ($A$) mAP
@@
71.74
Table 6: Effect of the components in computing the attention weights via Eq. 7
Table 7: Design choices of the relations and adaptive weights.
(a) Effect of the spatial and geometric relations.
Method mAP
$dx$ $dy$ $da$
70.84
   + dist
70.96
    + IoU
71.51
        + area
71.74
(b) Effect of the factors in computing $\beta$.
Attention Weights ($\beta$) mAP
@@
- scale ($\epsilon=\sqrt{S}=32$) 70.35(+3.49)
- overlap ($\delta=1$) 67.32(+0.46)
- density ($\bar{\rho}_{i}=0$) 69.85(+2.99)
Ours 71.74(+4.88)
4.4 Ablation Study
To understand the effect of each design choice in our model, we conduct a series of experiments to shed light on the following:
4.4.1 Preliminary Detection Phase
Compared to the standard detection pipeline, our model incorporates two detection heads - placed before and after the Transformer module. The output from the initial detection phase, termed ’Preliminary Detections’, includes a classification result (parameterized as a softmax distribution) from the first head, which forms a component of the RoI token. We posit that knowing the class information with uncertainties would help with reasoning about the inter-object relationships. To empirically validate this hypothesis, we compared the performance of our model with and without training the first detection head. As Tab. 6 shows, although solely incorporating the Transformer offers an improvement of mAP to the baseline, omitting the preliminary detection leads to a notable decline in performance. This suggests that relying only on the Transformer for RoIs to interact lacks efficacy. In contrast, the explicit inclusion of preliminary classification data, despite its potential inaccuracies, enhances the model’s ability to reason about semantical and contextual relationships. The results underscore the value of early classification cues in guiding the relational reasoning process within our proposed architecture.
4.4.2 Spatial and Geometric Relations
The different terms presented in Tab. 1 characterize various aspects of the spatial and geometric relationships among objects (RoI Tokens) within an image. In this section, we aim to empirically evaluate the individual and collective contributions of these spatial and geometric relational terms to the overall performance of our detection model. As shown in Tab. 7(a), IoU and rel. area contribute the most. Intuitively, they are particularly helpful when reasoning about the co-occurrence and spatial arrangement of objects. For example, IoU helps to disambiguate the overlapping, potentially duplicate or conflicting detections. Similarly, relative area aids in discerning the size relationship between objects. Consequently, our method can effectively solve the problem in the motivation example. See Sec. 5.2 for details.
4.4.3 Adaptive Attention Weights
As mentioned in Sec. 3.3, making the attention weights adaptive to specific RoI Tokens is essential to cope with the diversity and complexities in a scene. We evaluate our density- and scale-aware attention weighting scheme which is designed to augment the scaled-dot-product self-attention and allow the model to dynamically adjust its focus based on the scale of objects and their surrounding density. Findings in Tab. 7(b) indicate that masking the influence of overlapping RoIs plays a crucial role. This observation aligns with our initial understanding that indiscriminately emphasizing neighboring RoIs, without considering overlap, could lead to skewed attention distributions and potentially impair the model’s ability to accurately discern between distinct objects.
@@ 5 Analysis
5.1 Evidential Statistics
@@
$d(S_{1},S_{2})=\frac{1}{2}(\frac{1}{|S_{1}|}\sum_{i\in S_{1}}\min_{j\in S_{2}}\|\text{dist}_{ij}\|^{2}_{2}+\frac{1}{|S_{2}|}\sum_{j\in S_{2}}\min_{i\in S_{1}}\|\text{dist}_{ij}\|^{2}_{2}).$ (8)
@@
5.2 Qualitative Comparison
Figure 5: Qualitative comparison of our model and ReDet. Our method effectively reduces semantically incompatible detections, e.g., Small Vehicle and SHip on tennis courts.
5.2.1 Conflict Removal
Perhaps the most obvious advantage of considering inter-object relationships is that it helps to ensure the arrangement of detected objects aligns with real-world expectations and dataset distribution. For example, a basketball or tennis court should be no larger than an airplane. Two airplanes in the same image should have consistent sizes. And most importantly, the nose of a plane, despite the resemblance in some way, should not be classified as a ship. These examples (Fig. 1, Fig. 5) underscore the model’s proficiency in removing conflicts and ensuring logical and physically feasible detections.
5.2.2 Co-occurrence Recognition
Beyond resolving conflicts, our model exhibits a remarkable ability to leverage patterns of co-occurrence for enhanced detection accuracy. An illustrative case is the identification of a fleet of aligned planes (Fig. 5, third column). This aspect of co-occurrence pattern recognition is particularly beneficial in complex aerial images where objects often appear in structured groups or formations.
5.2.3 Failure Case
Unfortunately, there are cases where incorrect understanding of relationships could lead to even worse results, as depicted in Fig. 6.
Figure 6: An example failure case. The harbor and ship detections are all incorrect but reinforce each other.
@@ 6 Conclusion
In conclusion, this study presents a novel approach to object detection in aerial imagery, leveraging a Transformer-based architecture augmented with a two-stage detection process and adaptive attention mechanisms. Our model effectively captures and utilizes the spatial and geometric relationships among objects, as evidenced by its improved performance in handling diverse and complex scenes. The introduction of preliminary detection heads and the innovative use of scale- and density-aware attention weighting schemes have been shown to be particularly effective in enhancing detection accuracy.
Despite its strengths, our model has certain limitations. Firstly, the inclusion of the Transformer module notably increases computational demands, although the impact on training time is less pronounced. Secondly, inaccuracies in understanding object relationships can occasionally degrade detection results, as highlighted in some failure cases. Lastly, the model’s design incorporates several choices that may be specific to the dataset and application domain. Future efforts should aim to mitigate these limitations, striving towards a more universally applicable framework.

### v2

@@ 4.4 Ablation Study
Self supervised loss w/o 71.89
w 72.16
Multi-level fusion w/o 71.17
w 72.16
Table 5: Ablation of detection performance w/ or w/o Preliminary detection training, Self-supervised loss, Multi-level fusion.
RoI Tokens CLIP Tokens Relations Ada. Weight mAP50
@@
71.61
72.16
Table 6: Effect of the components in computing the attention weights via Eq. 6. Relations are calculated in Eq. 2, and Adaptive weight is calculated in Eq. 3
4.5 Self-supervised and Multi-level Fusion
We show the additional improvements achieved by incorporating the self-supervised loss and multi-level fusion for CLIP tokens in Tab. 5. Self-supervised loss effectively regularizes CLIP tokens to prevent representation collapse. And multi-level fusion helps capture information at different spatial scales.
$dx$ $dy$ $d\alpha$ dist IoU area mAP50
70.79
71.03
71.46
72.16
Table 7: Effect of the spatial and geometric relations.
Attention Weights ($\beta$) mAP50
@@ 4.5 Self-supervised and Multi-level Fusion
- scale ($\epsilon=\sqrt{S}=32$) 70.49
- density ($\bar{\rho}_{i}=0$) 71.30
Ours 72.16
Table 8: Effect of the factors in computing $\beta$.
@@ 5 Analysis
Figure 3: Qualitative comparison of our model and ReDet.
5.1 Evaluation Statistics
@@
$\begin{split}d(S_{1},S_{2})&=\frac{1}{2}\left(\frac{1}{|S_{1}|}\sum_{i\in S_{1}}\min_{j\in S_{2}}\|\text{dist}_{ij}\|^{2}_{2}\right)\\ &\quad+\frac{1}{2}\left(\frac{1}{|S_{2}|}\sum_{j\in S_{2}}\min_{i\in S_{1}}\|\text{dist}_{ij}\|^{2}_{2}\right)\end{split}$ (8)
@@
5.2 Limitations
While our method effectively captures inter-object relationships to improve detection accuracy, there are cases where this approach can lead to undesirable results. Specifically, when a wrong detection occurs for one object, it can propagate errors to nearby objects, particularly if those objects share similar spatial or semantic characteristics. As shown in Fig. 5, the false detection of ships in the ReDet model leads to an increased number of false positives for other ship instances when inter-object relationships are captured. This demonstrates that while our method enhances the overall detection performance, incorrect understanding or misidentification of one object may negatively influence the detection of surrounding objects, especially in cases where the objects are spatially or contextually similar.
Figure 5: An example of a failure case: The ship detections are all incorrect, but they reinforce each other, leading to an increased number of false ship detections.
@@ 6 Conclusion
In this work, we propose a Transformer-based framework to enhance object detection by effectively capturing inter-object and object-background relationships. By integrating the strengths of Transformer models with the cross-modal capabilities of CLIP, our approach not only improves the interaction between Region of Interest (RoI) proposals but also leverages background context for more accurate detections. Extensive experiments on several benchmark datasets demonstrate the effectiveness of our method, yielding consistent improvements over existing detectors. Our analysis further shows that the model reduces scale inconsistency and improves spatial and geometric understanding.